OpenStudy (jpsmarinho):

How to solve it Lim {(x^8 + x^4 )^(1/2)}/x^2 when x goes to 0. Is much hard with L'Hôspital , I think.

OpenStudy (amistre64):

factor out the top

OpenStudy (amistre64):

(x^8 + x^4 )^(1/2) = [x^4(x^4+1)]^(1/2) = x^2(x^4+1)^(1/2)

OpenStudy (amistre64):

as such, we have a removable discontinuity and we are left with: sqrt(x^4+1); at x=0

OpenStudy (jpsmarinho):

O_O thank God, you get it! Thanks!!